OpenStudy (anonymous):

by writing z in the form z=a+bi, find all solutions z of the following equations: 1) z^2=2+i 2) z^2-(3+i)z+(2+2i)=0

OpenStudy (anonymous):

simplifying gives\[a^2-b^2+2abi=2+i\]\[a^2-b^2-3a+b+2+(2ab-a-3b+2)i=0\]

OpenStudy (anonymous):

for example from first one \[a^2-b^2=2\]\[2ab=1\]combine 2 equations\[a^2-\frac{1}{4a^2}=2\]let \[t=a^2\]and solve

OpenStudy (anonymous):

I do not understand..... plz show me step by step.

OpenStudy (anonymous):

a^2-b^2+2abi=2+i a^2-b^2= 2 , 2abi=i => 2ab=1 I got this part, where did you get a^2-(1/4a^2)=2?

OpenStudy (anonymous):

for the first one you are solving \(z^2=2+i\) for \(z\) meaning you need both square root sof \(2+i\) if you are familiar with writing complex numbers in polar form, this would not be that hard. otherwise you have \((a+bi)^2=a^2-b^2+2abi\) and solve \[a^2-b^2=2, 2ab=1\] if you got that far, then \(2ab=1\implies b=\frac{1}{2a}\) making the first equation \[a^2-\frac{1}{4a^2}=2\]

OpenStudy (anonymous):

so for the first question. z=2. is that an answer?

OpenStudy (anonymous):

this kind of sucks because you are going to have to solve \[x^4-2x^2-\frac{1}{4}=0\]

OpenStudy (anonymous):

no 2 is not and answer because \(2^2=4\) and not \(2+i\)

OpenStudy (anonymous):

z1= (2+sqrt5)/2 , z2= (2-sqrt5)/2

OpenStudy (anonymous):

@satellite73 is that a right answer?